OpenStudy (anonymous):

Solve: 4(x + 1) + 8 = 24

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

4x + 4 +8 = 24 4x = 24 - 12 4x = 12 x = 3

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

x = 3.

OpenStudy (anonymous):

x=3